OpenStudy (anonymous):

Determine the area of the inner loop of the polar curve r = 1 − 2 sin (θ).

OpenStudy (anonymous):

Is this the right setup?

OpenStudy (anonymous):

Where r_1=1-2sin(theta)

OpenStudy (anonymous):

@jim_thompson5910 or @Abhisar please help

OpenStudy (freckles):

So we need to find when r=0 when theta is between 0 and 2pi So you have 1/2=sin(theta) which is when you said theta=5pi/6 or pi/6 So I think your limits are right

OpenStudy (tkhunny):

I get \(\pi - \dfrac{3}{2}\sqrt{3}\). Why do you doubt?

OpenStudy (anonymous):

I was just checking to make sure @tkhunny